OpenStudy (roberts.spurs19):

Please Please help! Inside the earth it is only the mass in the sphere beneath you that contributes to the gravitational pull on you. Assume that the Earth is a sphere of uniform density. What is the gravitational field strength halfway to the centre of the Earth? And three-quarters of the way to the centre of the Earth? Write down the general rule for the variation of g with distance from the centre of the Earth to its surface. Thank you!

OpenStudy (anonymous):

\[F _{g} = \frac{ Gm }{ r ^{2}}\] m = 5.97*10^24 kg r = 6.37*10^6m G = 6.67*10^-11Nm^2Kg^-2 now, obviously r changes, depending on where abouts you want to measure it from :)

OpenStudy (anonymous):

Hope it helps

OpenStudy (roberts.spurs19):

Thank you so much!!!!

OpenStudy (anonymous):

algood bro :)